candygirl220:

Divide me by 7, the remainder is 5. Divide me by 3, the remainder is 1 and my quotient is 2 less than 3 times my previous quotient. What number am I?

Vocaloid:

let's let the mystery number be y the first quotient can be x the second quotient can be z y/7 = x + 5 y/ 3 = z + 1 "second quotient is 2 less than 3 times my previous quotient" z = 3x - 2 plug this into z into the second equation to make both equations in terms of x and y and then solve by setting up a system of equations and using elimination to find y

Vocaloid:

sorry if this is taking a while I'm on mod duty and I'm dealing with trolls

candygirl220:

its okay

Vocaloid:

yeah I got 168 too gj

candygirl220:

thank you!

Vocaloid:

hm now that I'm double checking it it doesn't quite work let me redo it ;_;

candygirl220:

ok

Vocaloid:

sorry I'm really drawing a blank here as to how to do this systematically @563blackghost @sillybilly123

563blackghost:

Are you finding y specifically or each solution? There three since there are three variables.

Vocaloid:

it's just asking for the initial number y so just y

563blackghost:

well I got 168 as well O.O

Vocaloid:

yeah the problem is that 168 doesn't satisfy the first condition (it's divisible by 7 so no remainder) ;_; I don't know how to set up the equation so that x and z are integers only

563blackghost:

hmmm

Vocaloid:

I have a friend who knows a bit of discrete math, I'm going to ask if she can help ;; you can solve this using modulus operators

candygirl220:

ok

dude:

Possible numbers, 19 or 40

Vocaloid:

40 works, I don't think 19 satisfies the second condition though

dude:

Yep 13 = 3*5 - 2

sillybilly123:

\(n \equiv_7 5 \implies n = 7 \alpha + 5\) \(n \equiv_3 1 \implies n = 3 \beta + 1\) For the weirder bit ...."and my quotient is 2 less than 3 times my previous quotient".... I guess that means \(\beta = 3 \alpha - 2\) Which means that: \( 7 \alpha + 5 = 3 (3 \alpha - 2) + 1\) Or \(\alpha = 5\) so \(n = 40\)

candygirl220:

thank you everyone who helped!